OpenStudy (van.gosh):

A pulley system is used to lift a crate weighing 150 newtons straight up. If the force applied is 87 newtons and the ideal mechanical advantage of the pulley system is 2, what is the percentage efficiency of the pulley system?

OpenStudy (mathmate):

Hints: Effort (theoretical) = Force / mechanical advantage Effort (applied) = actual effort required. Efficiency = theoretical / actual.

OpenStudy (osprey):

@van.Gosh I see you could be a fan of Vincent, and maybe Rembrandt ?

OpenStudy (van.gosh):

@osprey I actually prefer the rococo style in general. :P

OpenStudy (van.gosh):

I'm still stuck on the problem. I'm not sure how all of these terms relate to each other very well. What am I supposed to do with the IMA?? :(

OpenStudy (osprey):

@IrishBoy123 Any thoughts for van.Gosh ?

OpenStudy (irishboy123):

Hi @osprey Following @mathmate 's hints in more concise terms.... Hints: Effort (theoretical) = Force / mechanical advantage \(\color{blue} { E_T = \dfrac{F}{MA} = \dfrac{150}{2} = 75N}\) Effort (applied) = actual effort required. \(\color{blue}{E_A = 87N}\) Efficiency = theoretical / actual. efficiency \(\color{blue}{\eta = \dfrac{75}{ 87}} \qquad \star\) That makes loadsa sense too because, if we should only need 75N to lift this mass with an idealised frictionless pulley, but we actually need 87N, then it makes sense to say that: \(\eta = \dfrac{75}{87}\) ....because it will be 100% efficient when the numerator and denominator are the same... Just to be clear, however ... if you are referring to my great artistic skills ( ;-| ), I much prefer cubism to any flavour of impressionism , though this is pretty awesome IMHO: 

OpenStudy (osprey):

@IrishBoy123 Yup that's some painting/lithographs too. Edvard Munch in what seems to be a far from happy mood. If I hadn't once pinned that up on an office wall somewhere, I'd be laughing even more hysterically than I am now. Nice one, and the artwork does you proud (even if you DIDN'T paint it !)

OpenStudy (blazeryder):

I got to go to the Van Gogh museum in Amsterdam this past summer. I am not a big fan of art, but I actually like his work!